OpenStudy (anonymous):

A study of two hundred teens found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 12 hours. The population standard deviation is 4 hours. What is the margin of error for a 98% confidence interval? 0.122 0.659 0.295 0.313

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kropot72):

The margin of error is given by: \[z \times \frac{\sigma}{\sqrt{n}}\] Do you know the value of z for a 98% confidence interval?

OpenStudy (anonymous):

I have the table but don't know how to find the z value.?

OpenStudy (kropot72):

It depends on the format of the table. Using my table I found the z value to be approximately 2.327. Now try using you table to find the cumulative probability for z = 2.327 and post the result.

OpenStudy (kropot72):

@mathhelpnow Are you there?

OpenStudy (anonymous):

Yes, just don't see 2.327 on my z-table. Outside, there are numbers on the outer left and top. Do I match up any numbers there?

OpenStudy (kropot72):

In the left hand column for z, find 2.3. Then move across to the left until you find .4901. This should be in the column headed .03 (or possibly just 3). Do this make sense?

OpenStudy (kropot72):

Does this make*

OpenStudy (anonymous):

Sorry, I'm new to the z-table; these are the numbers next to 2.3: 2.3 0.9893 0.9896 0.9898 0.9901 0.9904 0.9906 0.9909 0.9911 0.9913 0.9916

OpenStudy (kropot72):

We need to interpolate to find a z value for 0.49 (which is half of 0.98 or 98%). In any case, using my estimate for z, we can find the margin of error from: \[2.327\times\frac{4}{\sqrt{200}}=you\ can\ calculate\]

OpenStudy (kropot72):

Using your format of normal distribution table, you need to estimate the z value for .9900.

OpenStudy (kropot72):

I appears that your table has negative values of z as well as positive values.

OpenStudy (anonymous):

Ok, How do I estimate that? 0.313 because it would be under the 0.3 column?

OpenStudy (kropot72):

In your table, a z value of 2.33 gives a cumulative probability of .9901. Does that make sense?

OpenStudy (anonymous):

Yes,,, so I use .9901?

OpenStudy (kropot72):

Looking at the answer choices, it appears that a z value of 2.33 can be used in the calculation.

OpenStudy (kropot72):

So the margin of error is: \[2.33\times\frac{4}{\sqrt{200}}=you\ can\ calculate\]

OpenStudy (anonymous):

Does the square root 200 make it: 2.33 x 4 / 10 sqrt 2 ?

OpenStudy (kropot72):

Yes: \[\frac{2.33\times4}{10\times \sqrt{2}}\]

OpenStudy (anonymous):

Okay. 9.32 / 14.14 = 0.659!

OpenStudy (kropot72):

Correct. Good work!

OpenStudy (anonymous):

yay, thanks for the help:)

OpenStudy (kropot72):

You're welcome :)